OpenStudy (anonymous):

cot x sec4x = cot x + 2 tan x + tan3x

OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation.

OpenStudy (michele_laino):

I think that we have to use these identities: \[\Large \begin{gathered} {\left( {\sec x} \right)^4} = \frac{1}{{{{\left( {\cos x} \right)}^4}}} \hfill \\ \hfill \\ {\left( {\tan x} \right)^3} = {\left( {\frac{{\sin x}}{{\cos x}}} \right)^3} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

so, left side is: \[\Large \cot x{\left( {\sec x} \right)^4} = \frac{{\cos x}}{{\sin x}}\frac{1}{{{{\left( {\cos x} \right)}^4}}} = ...?\] please simplify

OpenStudy (anonymous):

\[\cos x/\sin x \times \left( cosx \right)^4\]

OpenStudy (michele_laino):

please you have to simplify cos(x) with (cos x)^4

OpenStudy (anonymous):

1/sin x * (cosx)^3

OpenStudy (michele_laino):

ok! correct!

OpenStudy (michele_laino):

now, we have to make the right side, equal to your expression for left side

OpenStudy (michele_laino):

so, we can write: \[\large \cot x + 2\tan x + {\left( {\tan x} \right)^3} = \frac{{\cos x}}{{\sin x}} + 2\frac{{\sin x}}{{\cos x}} + {\left( {\frac{{\sin x}}{{\cos x}}} \right)^3}\]

OpenStudy (michele_laino):

or: \[\large \begin{gathered} \cot x + 2\tan x + {\left( {\tan x} \right)^3} = \frac{{\cos x}}{{\sin x}} + 2\frac{{\sin x}}{{\cos x}} + {\left( {\frac{{\sin x}}{{\cos x}}} \right)^3} = \hfill \\ \hfill \\ = \frac{{\cos x}}{{\sin x}} + 2\frac{{\sin x}}{{\cos x}} + \frac{{{{\left( {\sin x} \right)}^3}}}{{{{\left( {\cos x} \right)}^3}}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

what is the common denominator?

OpenStudy (anonymous):

sinx*cosx^3

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

so, we can write this: \[\large \begin{gathered} \cot x + 2\tan x + {\left( {\tan x} \right)^3} = \frac{{\cos x}}{{\sin x}} + 2\frac{{\sin x}}{{\cos x}} + {\left( {\frac{{\sin x}}{{\cos x}}} \right)^3} = \hfill \\ \hfill \\ = \frac{{\cos x}}{{\sin x}} + 2\frac{{\sin x}}{{\cos x}} + \frac{{{{\left( {\sin x} \right)}^3}}}{{{{\left( {\cos x} \right)}^3}}} = \hfill \\ \hfill \\ = \frac{{\cos x{{\left( {\cos x} \right)}^3} + 2\sin x\sin x{{\left( {\cos x} \right)}^2} + \sin x{{\left( {\sin x} \right)}^3}}}{{\sin x{{\left( {\cos x} \right)}^3}}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ok I get it

OpenStudy (michele_laino):

we can simplify like this: \[\large \begin{gathered} = \frac{{\cos x}}{{\sin x}} + 2\frac{{\sin x}}{{\cos x}} + \frac{{{{\left( {\sin x} \right)}^3}}}{{{{\left( {\cos x} \right)}^3}}} = \hfill \\ \hfill \\ = \frac{{\cos x{{\left( {\cos x} \right)}^3} + 2\sin x\sin x{{\left( {\cos x} \right)}^2} + \sin x{{\left( {\sin x} \right)}^3}}}{{\sin x{{\left( {\cos x} \right)}^3}}} = \hfill \\ \hfill \\ = \frac{{{{\left( {\cos x} \right)}^4} + 2{{\left( {\sin x} \right)}^2}{{\left( {\cos x} \right)}^2} + {{\left( {\sin x} \right)}^4}}}{{\sin x{{\left( {\cos x} \right)}^3}}} = \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

\[(\cos(x)^4 +2\sin(x)^2(cosx)^2+\sin(x)^4)/\sin(x)xcos(x)^3\]

OpenStudy (michele_laino):

hint: the numerator is a perfect square

OpenStudy (michele_laino):

hint: \[\Large {\left\{ {{{\left( {\sin x} \right)}^2} + {{\left( {\cos x} \right)}^2}} \right\}^2} = ...?\]

OpenStudy (anonymous):

(sinx)^4+2(sinx)^2(cosx)^2+(cosx)^2

OpenStudy (anonymous):

(cosx)^4

OpenStudy (anonymous):

((sinx^2 + cosx^2)^2)/sinx(cosx)^3

OpenStudy (michele_laino):

I got this: since \[\Large \begin{gathered} {\left( {\cos x} \right)^4} + 2{\left( {\sin x} \right)^2}{\left( {\cos x} \right)^2} + \left( {\sin x} \right) = \hfill \\ = {\left\{ {{{\left( {\sin x} \right)}^2} + {{\left( {\cos x} \right)}^2}} \right\}^2} = {1^2} = 1 \hfill \\ \end{gathered} \] we have: \[\large \begin{gathered} \cot x + 2\tan x + {\left( {\tan x} \right)^3} = \frac{{\cos x}}{{\sin x}} + 2\frac{{\sin x}}{{\cos x}} + {\left( {\frac{{\sin x}}{{\cos x}}} \right)^3} = \hfill \\ \hfill \\ = \frac{{\cos x}}{{\sin x}} + 2\frac{{\sin x}}{{\cos x}} + \frac{{{{\left( {\sin x} \right)}^3}}}{{{{\left( {\cos x} \right)}^3}}} = \hfill \\ \hfill \\ = \frac{{\cos x{{\left( {\cos x} \right)}^3} + 2\sin x\sin x{{\left( {\cos x} \right)}^2} + \sin x{{\left( {\sin x} \right)}^3}}}{{\sin x{{\left( {\cos x} \right)}^3}}} = \hfill \\ \hfill \\ = \frac{{{{\left( {\cos x} \right)}^4} + 2{{\left( {\sin x} \right)}^2}{{\left( {\cos x} \right)}^2} + {{\left( {\sin x} \right)}^4}}}{{\sin x{{\left( {\cos x} \right)}^3}}} = \hfill \\ \hfill \\ = \frac{1}{{\sin x{{\left( {\cos x} \right)}^3}}} \hfill \\ \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok I understand thanks

OpenStudy (michele_laino):

sorry, another typo... \[\Large \begin{gathered} {\left( {\cos x} \right)^4} + 2{\left( {\sin x} \right)^2}{\left( {\cos x} \right)^2} + {\left( {\sin x} \right)^4} = \hfill \\ = {\left\{ {{{\left( {\sin x} \right)}^2} + {{\left( {\cos x} \right)}^2}} \right\}^2} = {1^2} = 1 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ok thanks

OpenStudy (michele_laino):

:)